OpenStudy (anonymous):

I'm supposed to write the equation of the line, in standard form, that has an x-intercept of 2 and is parallel to 2x + y = -5. Can someone help me through this problem?

OpenStudy (texaschic101):

we need to find the slope of 2x + y = -5...because a parallel line will have the same slope. Do you know how to find the slope of that equation ?

OpenStudy (anonymous):

Yes you subtract y2-y1/x2-x1?

OpenStudy (texaschic101):

no....you can do it that way if you have 2 sets of points..but you do not. So we must put the equation in y = mx + b form and m will be your slope. Can you do that ?

OpenStudy (anonymous):

That's what I thought but how would I put it into y=mx+b form? Would I subtract 2x from both sides?

OpenStudy (texaschic101):

yes...that is exactly what you would do.

OpenStudy (anonymous):

Meaning that the slope would be 2?

OpenStudy (texaschic101):

not quite....your subtracting 2x from both sides 2x + y = -5 2x - 2x + y = -2x - 5 y = -2x - 5 meaning the slope is -2. understand so far ?

OpenStudy (anonymous):

Yeah I've got it

OpenStudy (texaschic101):

now we will use y = mx + b slope(m) = -2 (2,0)....x = 2 and y = 0 now we sub 0 = -2(2) + b 0 = -4 + b 4 = b so your parallel equation is : y = -2x + 4 but we need it in standard form of Ax + By = C... y = -2x + 4 2x + y = 4 check to see if x intercept is 2 by subbing in 0 for y 2x + 0 = 4 2x = 4 x = 2 (correct) so your parallel equation in standard form where your x intercept is 2 is : 2x + y = 4

OpenStudy (texaschic101):

if you have any questions, ask them now

OpenStudy (anonymous):

No, wow thank you, you explained it quite well :)